# Are You Getting What You Pay For? Auditing Model Substitution in LLM APIs

### v1

@@ Abstract
The proliferation of Large Language Models (LLMs) accessed via black-box APIs introduces a significant trust challenge: users pay for services based on advertised model capabilities (e.g., size, performance), but providers may covertly substitute the specified model with a cheaper, lower-quality alternative to reduce operational costs. This lack of transparency undermines fairness, erodes trust, and complicates reliable benchmarking. Detecting such substitutions is difficult due to the black-box nature, typically limiting interaction to input-output queries. This paper formalizes the problem of model substitution detection in LLM APIs. We systematically evaluate existing verification techniques, including output-based statistical tests, benchmark evaluations, and log probability analysis, under various realistic attack scenarios like model quantization, randomized substitution, and benchmark evasion. Our findings reveal the limitations of methods relying solely on text outputs, especially against subtle or adaptive attacks. While log probability analysis offers stronger guarantees when available, its accessibility is often limited. We conclude by discussing the potential of hardware-based solutions like Trusted Execution Environments (TEEs) as a pathway towards provable model integrity, highlighting the trade-offs between security, performance, and provider adoption. Code is available at https://github.com/sunblaze-ucb/llm-api-audit
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities, leading to their widespread adoption through cloud-based APIs (OpenAI, 2022; Anthropic, 2023; Google, 2023; TogetherAI, 2023). Users typically select and pay for services based on advertised model specifications, such as parameter count (e.g., Llama 405B vs 70B), training data, or benchmark performance (Touvron et al., 2023; Bai et al., 2023; Mistral, 2023). However, the operational complexity and substantial computational cost associated with hosting state-of-the-art LLMs create incentives for service providers to optimize resource usage. This optimization can sometimes manifest as model substitution, where the provider covertly uses a different model than the one advertised and billed to the user.
Such substitutions can occur for various reasons. A provider might replace a large, expensive model (e.g., Llama-3.1-405B) with a smaller, cheaper one (e.g., Llama-3.1-70B) or a heavily quantized version (e.g., INT8 or FP8) to reduce GPU memory and compute costs, thereby increasing profit margins (Gao et al., 2024; Sun et al., 2024b). Alternatively, substitutions might happen for operational reasons like load balancing; if the requested model variant is under heavy load, requests might be rerouted to a different, less loaded variant (e.g., a version fine-tuned on different data). While potentially benign in intent, undisclosed substitutions break the implicit contract with the user, compromise the reliability of services built upon these APIs, hinder reproducible research, and invalidate benchmarks performed through third-party APIs. Furthermore, subtle changes from quantization or fine-tuning might inadvertently affect model safety or introduce biases. Imagine a researcher relying on a specific model’s advertised capabilities for scientific analysis, only to unknowingly receive results from a substituted, less capable model, potentially invalidating their findings.
Detecting model substitution is challenging in a black-box setting. Users typically only have query access, receiving text outputs and sometimes limited metadata like token log probabilities (Carlini et al., 2024). They lack direct access to model weights or the underlying inference infrastructure. This information asymmetry makes it difficult to verify the provider’s claims. Dishonest providers might further employ active countermeasures, such as detecting and evading benchmark queries or randomly mixing outputs from the correct and substituted models to obscure the signal from auditing attempts.
This paper addresses the critical problem of auditing model substitution in black-box LLM APIs. We formalize the verification task and systematically investigate the efficacy and robustness of various detection methodologies under realistic adversarial scenarios. Our contributions are:
@@
  • We design and analyze several practical adversarial attack scenarios relevant to model substitution, including quantization, randomized substitution, benchmark evasion, and limiting information disclosure.
  • We empirically evaluate the effectiveness of existing and potential detection techniques (text classification, identity prompting, model equality testing, benchmark analysis, log probability verification, embedding subspace analysis) against these attacks.
  • We analyze the robustness of these methods, showing that output-based statistical tests like Maximum Mean Discrepancy (MMD) are resilient to certain attacks but vulnerable to others, while log probability analysis offers stronger guarantees when available.
  • We discuss the potential of hardware-based solutions like Trusted Execution Environments (TEEs) as a path towards provable integrity, albeit requiring provider adoption.
@@ 2 Related work
LLM API auditing and monitoring.
Several studies have monitored the behavior and performance of commercial LLM APIs over time. Chen et al. (2023) tracked changes in ChatGPT’s capabilities, highlighting behavioral drift. Eyuboglu et al. (2024) characterized updates to API-accessed ML models. Gao et al. (2024) specifically tested API output text distributions against reference distributions using Maximum Mean Discrepancy (MMD). We extend this by evaluating MMD and other methods against adversarial substitution attacks. Complementary work by Sam et al. (2025) predicts black-box LLM performance using self-queries, aligning with our goal of verifying model identity.
Detecting LLM-generated text.
Detection of LLM-generated text has been extensively studied, including post-hoc methods and proactive watermarking techniques (Yang et al., 2023; Ghosal et al., 2023). Zero-shot detection approaches use stylistic features or model-specific ”fingerprints” to distinguish AI-generated content without specialized training data (Mitchell et al., 2023; Bao et al., 2023; Yang et al., 2024). Trained classifiers (Hu et al., 2023; Sun et al., 2025) utilize datasets from various sources to differentiate model outputs. Although effective against stylistically distinct models, these approaches may struggle with subtle substitutions within the same model family. LLM watermarking techniques embed hidden signals in outputs to trace content ownership (Kirchenbauer et al., 2023; Zhao et al., 2024a; b), but they are provider-centric and not intended for end-user verification of model identity, particularly as the same watermark might span different backend models.
Verifiable computation for ML.
Cryptographic techniques like Zero-Knowledge Proofs (ZKPs) (Feng et al., 2021; Sun et al., 2024a; Xie et al., 2025) and hardware-based Trusted Execution Environments (TEEs) (NVIDIA, 2023) offer provable verification of ML inference. ZKPs allow proof of correct inference without revealing inputs or weights but face substantial overhead for large LLMs (Sun et al., 2024a). TEEs provide hardware-level guarantees with lower performance costs but depend heavily on specific infrastructure (NVIDIA, 2023). We highlight TEEs as promising yet infrastructure-dependent, focusing primarily on methods practical today with existing API structures.
@@ 3.1 Problem formulation
We consider a scenario involving three main entities:
  • User/Auditor: Wishes to use an LLM service for a task involving input prompts $x$ sampled from a distribution $\pi(x)$.
  • Service provider: Offers access to an LLM via a black-box API. The provider advertises that the API uses a specific target model $M_{spec}$ (e.g., Llama-3-70B-Instruct).
  • Target model ($M_{spec}$): The model the user expects and pays for, characterized by a conditional probability distribution $P_{spec}(y|x)$ over output sequences $y$.
The core problem arises because the provider might, intentionally or unintentionally, use an alternative model $M_{alt}$ while still claiming to use $M_{spec}$. The alternative model $M_{alt}$ could be: (1) A smaller model from the same family (e.g., 8B instead of 70B). (2) A quantized version of $M_{spec}$ (e.g., INT8 or FP8). (3) A model from an entirely different family (e.g., Mistral instead of Llama) (4) A fine-tuned variant of $M_{spec}$, etc.
The user interacts with the API by sending a prompt $x$ and receiving an output $y$ generated according to the provider’s actual backend distribution $P_{actual}(y|x)$. Depending on the API, the user might also receive additional information, such as log probabilities of the generated tokens. The user typically has reference access (e.g., via local execution or a trusted provider) to the target distribution $P_{spec}(y|x)$ and possibly some alternative distributions $P_{alt}(y|x)$.
Auditing goal.
The user/auditor aims to determine, with high confidence, whether the provider is faithfully using the specified model. Formally, given samples $(x,y)$ from the provider’s API ($y\sim P_{actual}(\cdot|x)$) and potentially reference samples from $P_{spec}(\cdot|x)$, the auditor wants to test the null hypothesis:
[…]
@@
[…]
An effective verification method should achieve:
  • High accuracy: Low False Positive Rate (incorrectly flagging an honest provider) and low False Negative Rate (failing to detect a substitution).
  • Efficiency: Require a reasonable number of queries and computational overhead.
  • Robustness: Remain effective even when the provider employs countermeasures.
  • Generality: Apply across different models, tasks (prompts $\pi$), and levels of API access.
@@ 3.2 Adversarial attack scenarios
A dishonest provider aiming to perform model substitution while avoiding detection might employ several evasion strategies. We analyze the following practical attack scenarios:
Quantization substitution.
The provider replaces the full-precision target model $M_{spec}$ with a quantized version (e.g., INT8, FP8, NF4). Quantization significantly reduces memory footprint and often accelerates inference, lowering costs. While preserving much of the model’s capabilities, it slightly alters the output distribution $P_{alt}(y|x)$. The attack relies on this distributional difference being too small for naive detection methods to pick up, especially with limited samples.
Randomized model substitution.
To make detection harder, especially if simpler substitutions prove detectable, the provider can employ randomized substitution. A query is routed to the original model $M_{spec}$ with probability $p$ and to a cheaper substitute $M_{alt}$ with probability $1-p$. The resulting output distribution observed by the user is a mixture:
$P_{mixed}(y|x)=p\cdot P_{spec}(y|x)+(1-p)\cdot P_{alt}(y|x).$
As $p$ approaches 1 (low substitution rate), $P_{mixed}$ becomes statistically closer to $P_{spec}$, making detection via sampling-based tests increasingly difficult. A sophisticated provider might adaptively adjust $p$, perhaps increasing it (reducing substitution) if query patterns suggest an audit (e.g., repeated prompts, known benchmark queries), aiming to balance cost savings with evasion.
Benchmark evasion (cached/routed output).
This attack specifically targets verification methods relying on known, fixed prompts, such as benchmark datasets or simple identity queries (”What model are you?”). The provider implements a detection mechanism (e.g., prompt hashing, string matching, embedding similarity with known benchmark prompts) to identify likely audit queries. Upon detection, the provider serves a response generated by (or routes the query to) a genuine $M_{spec}$ instance. If necessary, corresponding metadata like log probabilities can also be cached or generated on-demand from $M_{spec}$. For all other ”normal” user queries, the cheaper $M_{alt}$ is used. This allows the provider to pass audits based on known inputs while still achieving cost savings on the majority of traffic.
Limiting information disclosure.
Providers might proactively limit the information exposed through their APIs to thwart known verification techniques that rely on richer metadata. For instance, after researchers demonstrated that full logit distributions could be reconstructed using logit biasing parameters (Carlini et al., 2024; Finlayson et al., 2024), several major API providers removed or restricted this feature.
@@ 4 Model verification techniques and robustness analysis
Anyscale Yes Full Control Top 5
Together.ai Yes Full Control Single Token
@@
AWS Bedrock Yes Full Control No
Nebius AI Yes Full Control No
@@
DeepSeek Yes Partial Control Top 20
OpenAI No Partial Control Top 20
Cohere No Full Control Single Token
Anthropic No Partial Control No
Table 1: Transparency and control across LLM API providers (March 2025). “Full Control” implies typical parameters like temperature, top-p, top-k, etc.
LLM service providers offer a wide range of interfaces, from basic text-only web chats to advanced APIs that allow control over decoding parameters and access to token log probabilities (Table 1 and Table 5). This variation determines which verification techniques can be applied.
We evaluate each technique based on its required access level: (1) text-only output, (2) access to additional information (e.g., log probabilities) or control via APIs, and (3) reliance on provider-supplied information (e.g., TEEs discussed in Section 4.3). For each method, we assess its effectiveness and robustness against the attacks described in Section 3.2.
@@ 4.1 Text-output-based verification
These methods rely solely on the textual completions generated by the API, making them universally applicable but potentially less sensitive.
4.1.1 Text classifier (Sun et al., 2025)
Method.
This approach leverages the idea that different LLMs possess unique stylistic ”fingerprints”. Using text outputs from different source models, an auditor can construct a labeled dataset $D=\{(x_{i},z_{i})\}_{i=1}^{N}$, where each $x_{i}$ is a generated text sample and $z_{i}\in\{1,\ldots,K\}$ indicates the model that generated it. A classifier is then trained by fine-tuning an encoder LLM along with a classification head, formally to find:
$\theta^{*}=\arg\min_{\theta}\sum_{(x_{i},z_{i})\in D}\mathcal{L}\bigl{(}M_{\theta}(x_{i}),z_{i}\bigr{)},$
where $\mathcal{L}$ denotes the cross-entropy loss. During auditing, generated text from the API is fed to the classifier to predict its source.
Attack: Quantization substitution.
We investigate if text classifiers can distinguish between a full-precision model and its quantized versions, a common cost-saving substitution. We generated text using Llama-3.1-70B, Gemma2-9B, Mistral-7B-v0.3, and Qwen2-72B, comparing their original versions against FP8 and INT8 quantized counterparts. Using prompts from Ultrachat (Ding et al., 2023) and sampling with temperature 0.6, we created datasets (10k train, 1k validation samples per model version). We trained binary classifiers using BERT (Devlin et al., 2019), T5 (Raffel et al., 2020), GPT-2 (Radford et al., 2019), and LLM2Vec (BehnamGhader et al., 2024) embeddings.

### v2

@@ Abstract
Commercial Large Language Model (LLM) APIs create a fundamental trust problem: users pay for specific models but have no guarantee that providers deliver them faithfully. Providers may covertly substitute cheaper alternatives (e.g., quantized versions, smaller models) to reduce costs while maintaining advertised pricing. We formalize this model substitution problem and systematically evaluate detection methods under realistic adversarial conditions. Our empirical analysis reveals that software-only methods are fundamentally unreliable: statistical tests on text outputs are query-intensive and fail against subtle substitutions, while methods using log probabilities are defeated by inherent inference nondeterminism in production environments. We argue that this verification gap can be more effectively closed with hardware-level security. We propose and evaluate the use of Trusted Execution Environments (TEEs) as one practical and robust solution. Our findings demonstrate that TEEs can provide provable cryptographic guarantees of model integrity with only a modest performance overhead, offering a clear and actionable path to ensure users get what they pay for. Code is available at https://github.com/sunblaze-ucb/llm-api-audit
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities, leading to widespread adoption through cloud-based APIs (OpenAI, 2022; Anthropic, 2023; Google, 2023; TogetherAI, 2023). Enterprises and researchers select and pay for specific models (e.g., Llama 405B vs 70B) based on advertised capabilities, performance on leaderboards, and expected behavior (Touvron et al., 2023; Bai et al., 2023; Mistral, 2023). However, this black-box access model operates on an implicit assumption of trust: that the provider will faithfully serve the requested model. The immense computational cost of hosting state-of-the-art LLMs creates a powerful economic incentive for providers to violate this trust through model substitution—covertly replacing the advertised model with a cheaper, less powerful alternative.
This substitution threat is not hypothetical. A provider might swap a flagship model for a smaller variant or a heavily quantized version (e.g., INT8/FP8) to reduce GPU costs and increase profit margins (Gao et al., 2024; Sun et al., 2024b). Substitutions can also occur for operational reasons, such as rerouting traffic from an overloaded high-end model to a less-utilized, lower-tier one (e.g., a version fine-tuned on different data). While the intent may not always be malicious, any undisclosed substitution breaks the service agreement, compromises the reliability of dependent applications, hinders reproducible research, and invalidates the benchmark results. Furthermore, subtle changes from quantization or fine-tuning might inadvertently affect model safety or introduce biases (Hong et al., 2024). Imagine a researcher relying on a specific model’s advertised capabilities for scientific analysis, only to unknowingly receive results from a substituted, less capable model, potentially invalidating their findings.
For a user or auditor, verifying the model behind a black-box API is fundamentally challenging. The interaction is typically limited to sending prompts and receiving text (and optionally, token log probabilities (Carlini et al., 2024)), with no direct access to the model’s weights or the underlying infrastructure. This information asymmetry heavily favors the provider. A sophisticated provider can exploit this by employing active countermeasures, such as detecting and routing benchmark queries to the genuine model while serving a substitute to regular traffic, or by randomly mixing outputs from different models to obscure the statistical signal of the substitution.
This paper confronts the critical problem of auditing model substitution in black-box LLM APIs. We formalize the verification task, analyze the limitations of existing software-based methods under realistic adversarial conditions, and present a robust solution. Our central thesis is that the inherent ambiguity and nondeterminism of software-level signals make them insufficient for reliable auditing. Instead, we argue that hardware-backed attestation via Trusted Execution Environments (TEEs) is the only currently viable mechanism to achieve strong, efficient, and provable integrity for LLM APIs. Our contributions are:
@@
  • We design and analyze practical adversarial scenarios for model substitution, including quantization, randomized substitution, and benchmark evasion.
  • We empirically evaluate the effectiveness of existing detection techniques, showing that text-based statistical tests lack power against subtle or randomized substitutions, while metadata-based methods like log probability comparison are defeated by production-level inference nondeterminism.
  • We propose and evaluate TEEs as a practical and deployable solution, demonstrating that they uniquely provide provable model integrity in black-box APIs. We show that TEEs offer strong security guarantees with only a modest performance cost, establishing them as the most actionable path toward substitution-resistant LLM services.
@@ 2 Related work
LLM API monitoring and auditing. Several studies have monitored the behavior and performance of commercial LLM APIs over time. Chen et al. (2023) track changes in ChatGPT’s capabilities, highlighting behavioral drift, while Eyuboglu et al. (2024) characterizes updates to API-accessed ML models. Closer to our work, Gao et al. (2024) specifically test API output text distributions against reference distributions using Maximum Mean Discrepancy (MMD). Other approaches include direct auditing via identity-style prompting (Huang et al., 2025) and predicting black-box LLM performance using self-queries (Sam et al., 2025). When model internals are available, TopLoc (Ong et al., 2025) uses locality-sensitive hashing over intermediate activations to produce proofs of correct execution. While these methods provide a foundation for auditing, their robustness against determined adversaries and in realistic, non-deterministic production environments (He & Lab, 2025) remains an open question, which we address in this work.
Detecting LLM-generated text. Detection of LLM-generated text has been extensively studied, including post-hoc methods and proactive watermarking techniques (Yang et al., 2023; Ghosal et al., 2023). Zero-shot detection approaches use stylistic features or model-specific ”fingerprints” to distinguish AI-generated content without specialized training data (Mitchell et al., 2023; Bao et al., 2023; Yang et al., 2024). Trained classifiers (Hu et al., 2023; Sun et al., 2025) utilize datasets from various sources to differentiate model outputs. Although effective against stylistically distinct models, these approaches may struggle with subtle substitutions within the same model family. LLM watermarking techniques embed hidden signals in outputs to trace content ownership (Kirchenbauer et al., 2023; Zhao et al., 2024a; b), but they are provider-centric and not intended for end-user verification of model identity, particularly as the same watermark might span different backend models.
Verifiable computation for ML. Cryptographic and hardware-based techniques can provably verify ML inference. Zero-Knowledge Proofs (ZKPs) allow proof of correct inference without revealing inputs or model weights but face substantial computational costs for large LLMs, making them impractical for real-time APIs today (Sun et al., 2024a; Xie et al., 2025). In contrast, Trusted Execution Environments (TEEs) provide hardware-level guarantees of integrity and confidentiality with much lower overhead (NVIDIA, 2023). Our work builds on this insight, arguing that TEEs represent the most mature and practical path toward verifiable LLM inference in the current ecosystem.
@@ 3.1 Problem formulation
Consider an LLM API service with three entities:
  • User/Auditor: Use LLM services for tasks with input prompts $x$ drawn from a distribution $\pi(x)$.
  • Service provider: Offers access to an LLM via a black-box API. The provider claims to serve a specific target model $M_{\text{spec}}$ but may substitute it with $M_{\text{alt}}$.
  • Target model ($M_{\text{spec}}$): The advertised model with distribution $P_{\text{spec}}(y|x)$.
Model substitution occurs when the provider’s actual backend distribution $P_{\text{actual}}(y|x)$ differs from the advertised $P_{\text{spec}}(y|x)$ without user disclosure. Substitutions include: (1) smaller models (e.g., 7B vs 70B), (2) quantized variants (INT8 vs FP16) of $M_{\text{spec}}$, (3) fine-tuned or updated versions with altered training data or objectives, or (4) entirely different model families.
Open-source vs. proprietary access. The user’s ability to audit depends on the type of model and API. For open-source models, an auditor can run $M_{\text{spec}}$ locally to get reference outputs and log probabilities. For proprietary models (e.g., GPT-4, Claude-4), the auditor only has black-box access, creating a significant information asymmetry.
Auditing goal. The user/auditor aims to determine whether the provider is faithfully using the specified model. Formally, given samples $(x,y)$ from the provider’s API ($y\sim P_{\text{actual}}(\cdot|x)$) and potentially reference samples from $P_{\text{spec}}(\cdot|x)$, the auditor wants to test the null hypothesis:
$H_{0}:P_{\text{actual}}(y|x)=P_{\text{spec}}(y|x)\quad\forall x\sim\pi(x)\quad\text{(Honest Provider)}$
@@
$H_{1}:P_{\text{actual}}(y|x)\neq P_{\text{spec}}(y|x)\quad\text{for some }x\sim\pi(x)\quad\text{(Substitution Occurred).}$
An effective verification method should achieve: (1) Reliability - low false positive/negative rates, (2) Efficiency - reasonable query budgets, (3) Robustness - resistance to adversarial countermeasures, and (4) Generality - applicability across model types and API configurations.
@@ 3.2 Adversarial attack scenarios
Quantization substitution. The provider replaces the full-precision target model $M_{\text{spec}}$ with a quantized version (e.g., INT8, FP8, NF4). Quantization significantly reduces memory footprint and often accelerates inference, lowering costs. While preserving much of the model’s capabilities, it slightly alters the output distribution $P_{\text{alt}}(y|x)$. The attack relies on this distributional difference being too small for naive detection methods to pick up, especially with limited samples.
Randomized model substitution. To make detection harder, the provider can route a query to a cheaper substitute $M_{\text{alt}}$ with probability $p$ and to the specified model $M_{\text{spec}}$ with probability 1p1-p. The observed distribution is $P_{\text{mixed}}(y\mid x)=p\cdot P_{\text{alt}}(y\mid x)+(1-p)\cdot P_{\text{spec}}(y\mid x).$ As $p\to 1$ (low substitution rate), $P_{\text{mixed}}$ approaches $P_{\text{spec}}$, making sampling-based detection increasingly difficult. A sophisticated provider might adapt $p$ downward only when traffic appears non-audit-like.
Benchmark evasion (cached/routed output). This attack targets verification methods relying on known, fixed prompts, such as benchmark datasets or identity queries. The provider detects likely audit queries (e.g., via hashing or embedding similarity) and routes them to a genuine $M_{\text{spec}}$ instance or returns cached outputs and metadata. Ordinary traffic is served by $M_{\text{alt}}$.
Limiting information disclosure. Providers might proactively limit the information exposed via APIs (e.g., removing full logits, restricting top-$k$ log probabilities) after demonstrations that such features can leak sensitive details (Carlini et al., 2024; Finlayson et al., 2024).
@@ 4 Model verification techniques and robustness analysis
Anyscale Yes Full Control Yes
Together.ai Yes Full Control Yes
@@
AWS Bedrock Yes Full Control Yes
Nebius AI Yes Full Control Yes
@@
DeepSeek Yes Full Control Top 20
OpenAI No Partial Control Depends on model
Cohere No Full Control Top 1
Anthropic No Full Control No
Table 1: Transparency and control across LLM API providers (September 2025). “Full Control” implies typical parameters like temperature, top-p, top-k, etc. Log probability availability and limits vary by model/version.
LLM service providers offer a wide range of interfaces, from basic text-only web chats to advanced APIs that allow control over decoding parameters and access to token log probabilities (Table 1 and Table 5). The level of information disclosure significantly constrains which verification techniques can be applied. In this section, we systematically evaluate various auditing methods across different access levels: (1) text-only output, (2) output plus metadata (e.g., log probabilities, activations), and (3) provider-supplied integrity (e.g., TEEs in Section 4.3). For each technique, we detail the methodology, describe our experimental setup for evaluating its robustness against the adversarial scenarios outlined in Section 3.2, and analyze its effectiveness.
@@ 4.1 Text-output-based verification
4.1.1 Text classifier
Method. Sun et al. (2025) use a trained classifier to predict the source model of generated samples, exploiting stylistic “fingerprints” unique to each model. Given a dataset of model completion labeled with model, the classifier is trained to predict the label using cross-entropy loss.